OpenStudy (anonymous):

Find a · b. |a| = 9, |b| = sqrt(14), the angle between a and b is 45°

OpenStudy (anonymous):

i fi remember right the dot product of the two is a*b*cos(angle)

OpenStudy (anonymous):

i used that but got the wrong answer somehow.

OpenStudy (anonymous):

i did \[9(\sqrt{14}) \cos(45)\] and it gave me 23.81

OpenStudy (anonymous):

what is the right answer

OpenStudy (anonymous):

i haven't been given the right answer yet

OpenStudy (anonymous):

are us missing anything in the question because i get the same answer as u

OpenStudy (raden):

do u have choices ?

OpenStudy (anonymous):

i copied and pasted the exact question on here. I don't have any choices just 5 attempts

OpenStudy (anonymous):

I've used 4/5 of my attempts

OpenStudy (raden):

actually, ur answer is right. but, maybe the answer wants represent in sqrt form. in trygonometry, we knowed that the exact value of cos45 deg = 1/sqrt(2) or 1/2 * sqrt(2). so, ur calculating can be : a. b = |a|*|b|*cos45 = 9 * sqrt(14) * 1/sqrt(2) = 9 * sqrt(7)

OpenStudy (anonymous):

RadEn could be right i checked my text book and nothing is wrong with the equation

OpenStudy (raden):

yeah, the question is not right... but, the answer can be formed sqrt value not decimal, maybe :)

OpenStudy (raden):

*is not wrong, i mean :)

OpenStudy (anonymous):

thanks a lot. 9(sqrt(7) was the right answer!!

OpenStudy (raden):

welcome :)